OpenStudy (anonymous):

Solve the system of equations x = –6y + 2 and 1/2x - y = 1. What is the y-coordinate of the solution? A. -2/3 B. 0 C. -1/4 D. 1/4

hartnn (hartnn):

plug in x = –6y + 2 in 1/2x - y = 1 what u get ?

OpenStudy (anonymous):

1/2 (-6y+2) -y =1

OpenStudy (anonymous):

Is this correct?

hartnn (hartnn):

good! yes, correct :) now multiply (1/2) to each terms inside 1/2 * -6y =...? 1/2* 2 =... ?

OpenStudy (anonymous):

-3y + 1 -y =1

OpenStudy (anonymous):

Is this correct?

hartnn (hartnn):

correct! go on :)

OpenStudy (anonymous):

-4y = 0

OpenStudy (anonymous):

y equals zero?

hartnn (hartnn):

good work! y=0 is correct :)

hartnn (hartnn):

if and when you need 'x' , you can plug in y=0 in any of the equation, x= -6*0+2 x=2 :)

OpenStudy (anonymous):

Thanks a lot!!! :)

hartnn (hartnn):

welcome ^_^